# Lie symmetry group methods for differential equations

### v10

@@ II. $c_{0}=0$, $b_{0}\neq 0$: In this case $f$ and $g$ satisfy
The solution of this system leads to an integrable subclass of (157) admitting a two-dimensional non-abelian Lie point symmetry with commutation relation (136) and $\mathbf{v}_{2}$ having the form (137) ($b_{0}=1$)
@@
or as the solution of the separable equation
$zz^{\prime}-\frac{1}{2}z^{2}+\frac{2M}{k^{2}}=0$
obtained by selecting the invariant $z(r)=e^{r}R$ as the new dependent variable. Finally, second integration gives
@@
satisfying the boundary conditions $\lim_{x\to\infty}y(x)=6/25$ and $\lim_{x\to-\infty}y(x)=0$. In [1], it was shown that for the special wave speed $c=5/\sqrt{6}$, for which the Fisher’s equation passes the necessary condition to be of Painlevé type (we already discussed this constraint on page 4.11 as a condition of integrability by quadratures), can be reduced to the canonical form $r^{\prime\prime}=6r^{2}$ by a transformation of the both independent and dependent variables of the form $y(x)=f(x)r(s)$, $s=h(x)$ with $f$ and $h$ appropriately chosen.
@@ 5 Group-invariant solutions
$Lu=u_{t}-xu_{y}-x^{5}u_{xx}=\delta(t)\delta(x-x_{0})\delta(y).$
@@
$\begin{split}&\mathbf{p}_{t}=\partial_{t},\quad\mathbf{p}_{x}=\partial_{x},\quad\mathbf{p}_{y}=\partial_{y},\\ &\mathbf{k}_{xt}=t\partial_{x}+x\partial_{t},\quad\mathbf{k}_{yt}=t\partial_{y}+y\partial_{t},\quad\mathbf{j}_{xy}=y\partial_{x}-x\partial_{y},\\ &\mathbf{d}=t\partial_{t}+x\partial_{x}+y\partial_{y}-\frac{u}{2}\partial_{u},\end{split}$ (224)
@@
where $\Delta_{n}=\sum_{\mu=1}^{n}u_{x_{\mu}x_{\mu}}$ is the Laplace operator and $x=(x_{0},x_{1},\ldots,x_{n})\in\mathbb{R}^{1,n}$. The symmetry algebra is $(n+2)(n+3)/2$-dimensional conformal algebra of the conformal group $\operatorname{Conf}(\mathbb{R}^{1,n})$ with basis

### v11

@@ II. $c_{0}=0$, $b_{0}\neq 0$: In this case $f$ and $g$ satisfy
The solution of this system leads to an integrable subclass of (179) admitting a two-dimensional nonabelian Lie point symmetry with commutation relation (146) and $\mathbf{v}_{2}$ having the form (147) ($b_{0}=1$)
@@
or as an exact equation using the integrating factor $2e^{r}$. Finally, second integration gives
@@
satisfying the boundary conditions $\lim_{x\to\infty}y(x)=6/25$ and $\lim_{x\to-\infty}y(x)=0$. In [1], it was shown that for the special wave speed $c=5/\sqrt{6}$, for which the Fisher’s equation passes the necessary condition to be of Painlevé type (we already discussed this constraint on page 4.11 as a condition of integrability by quadratures), can be reduced to the canonical form $r^{\prime\prime}=6r^{2}$ by a transformation of the independent and dependent variables of the form $y(x)=f(x)r(s)$, $s=h(x)$ with $f$ and $h$ appropriately chosen.
@@ 5 Group-invariant solutions
$Lu=u_{t}+xu_{y}-x^{5}u_{xx}=\delta(t)\delta(x-x_{0})\delta(y).$
@@
$\begin{split}&\mathbf{p}_{t}=\partial_{t},\quad\mathbf{p}_{x}=\partial_{x},\quad\mathbf{p}_{y}=\partial_{y},\\ &\mathbf{k}_{xt}=t\partial_{x}+x\partial_{t},\quad\mathbf{k}_{yt}=t\partial_{y}+y\partial_{t},\quad\mathbf{j}_{xy}=y\partial_{x}-x\partial_{y},\\ &\mathbf{d}=t\partial_{t}+x\partial_{x}+y\partial_{y},\end{split}$ (246)
@@
where $\Delta_{n}=\sum_{\mu=1}^{n}u_{x_{\mu}x_{\mu}}$ is the Laplace operator on $\mathbb{R}^{n}$ and $x=(x_{0},x_{1},\ldots,x_{n})\in\mathbb{R}^{1,n}$. The symmetry algebra is $(n+2)(n+3)/2$-dimensional conformal algebra of the conformal group $\operatorname{Conf}(\mathbb{R}^{1,n})$ with basis